OpenStudy (anonymous):

Help! how do I solve this? I WILL FAN AND MEDAL Solve and check. r/6=5 2/3

OpenStudy (anonymous):

@haleyelizabeth2017 ,@DanJS and @TheSmartOne plz help!!!

OpenStudy (anonymous):

oops @DanJS

OpenStudy (anonymous):

@justuu

OpenStudy (anonymous):

A. 34 B. 30 2/3 C. 20 D. 17/18

OpenStudy (danjs):

change the 5 2/3 to 17 / 3

OpenStudy (danjs):

r / 6 = 17 / 3

OpenStudy (anonymous):

I will fan if I have not already.

OpenStudy (danjs):

cross multiply that, 3r = 6*17

OpenStudy (danjs):

3r = 102 divide by 3 r = 102/3 = ...

OpenStudy (anonymous):

I think it is A.

OpenStudy (anonymous):

Am I right?

OpenStudy (danjs):

if you arent confident about doing the r = 102 / 3 , you didnt understand the steps up to that did you?

OpenStudy (anonymous):

I kinda got it.

OpenStudy (danjs):

r= 102 / 3 = 34 try 34 for r in the original expression and see if it makes sense. 34/6 = 5 2/3 ?

OpenStudy (danjs):

5 2/3 is the same as 17 / 3 double top and bottom, you get 34/6